OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form.

OpenStudy (anonymous):

[y-5]=2[x+1] 2x-y+7=O

OpenStudy (lgbasallote):

use the point slope form \[y - y_1 = m(x-x_1)\] where m is the slope x1 and y1 are the x and y coordinated

OpenStudy (anonymous):

the answer is 2x-y+7=O